OpenStudy (anonymous):

ut of 1,100 discs tested, 13 are defective. Estimate the number of defective discs in a batch of 41,000. A. about 485 discs B. about 970 discs C. about 4,850 discs D. about 49 discs

OpenStudy (anonymous):

Out**

OpenStudy (anonymous):

A) 12*41000/1100

OpenStudy (anonymous):

Could U Guys Please Explain How U Got The Answer? (:

OpenStudy (anonymous):

Okay, so 13 out of 1100 are defective. So: 13/1100 = y/41000 Cross multiply to get: y = (13x41000)/1100 y = 484.54545454 Which rounds to 485

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Ok Thanks U Guys! (:

OpenStudy (anonymous):

yea